# Time-Optimal Flight with Safety Constraints and Data-driven Dynamics

### v1

@@ V-A Simulation
At test time, we select the best policy from the training phase and run 10 episodes (equivalent to 30 laps). The results reported in Table I are based on the average lap time over the 10 episodes. Additionally, we report the success rate (SR) by counting the number of episodes where the drone successfully finishes all 3 laps without any gate collisions [7]. After simulation, we select the best policy obtained in the residual simulator and deploy it in the real world. Note that while BEM and the residual simulator provide comparable accuracy levels, the latter is preferred due to its significantly lower computational demand - approximately one tenth that of the BEM simulator.
@@ V-A1 TuRBO vs WML
We first investigate the difference between WML and TuRBO, and discuss several advantages of TuRBO. We train the baseline MPCC [14] using both variants. Despite both methods achieving comparable maximum rewards, in this section we deive into understanding how each of these explore the parameter space. In Figure 4 we illustrate two of the tuned parameters - the minimum contour weight $Q_{c}$ and the progress weight $\mu$, across the number of episodes. WML starts a new trial every 100 iterations, while TuRBO runs multiple trust regions in parallel, facilitating a broader exploration of the parameter space. This is primarily due to TuRBO’s ability to select the local instances in an informed manner, as opposed to WML’s approach of random trial restarts. An additional advantage of TuRBO is that the trust regions are independent, allowing parallel execution without incurring extra computational cost. Consequently we advocate for TuRBO as our preferred tuning algorithm, for two reasons: i) comprehensive exploration of the entire parameter space; ii) computational efficiency and scalability, which are crucial for exploring large parameter spaces.
Figure 4: Parameter exploration using WML vs TuRBO. The different data samples are colored based on the collected reward. The figure shows that the WML method is prone to get stuck in local minima, while the TuRBO method is able to explore a broader parameter space.
V-A2 MPCC vs SMPCC
@@
We train both MPCC and SMPCC using TuRBO. We make some practical adjustments to the SMPCC formulation introduced in Section IV-A. Specifically, we relax the hard constraints of the tunnel and impose soft constraints instead, which is a standard practice in MPC and can be performed in a systematic way [64]. This has several benefits: i) improves the numerical stability of the solver; ii) handles infeasibilities which arise from model mismatch by allowing minor violations; iii) maintains the low computational complexity of MPCC. This adaption does not impose any drawbacks in terms of performance/safety. The soft constraint is implemented using a barrier function $p(h(x))$ to embed the constraint into the cost function. For nonlinear constraints defined as $h(x)\geq 0$, we employ:
$p(h(x))=\log\left(1+\exp\left(-\alpha h(x)\right)\right).$ (15)
@@
setting the penalty slope to $\alpha=100$. Our results, as detailed in Table I demonstrate SMPCC’s superiority over MPCC in terms of lap times and success rates. Due to spatial constraints (9), SMPCC can prevent crashes consistently without compromising the lap time. Allowing the controller to independently navigate within the tunnel, instead of following a predefined path, gives it greater flexibility to identify the best trajectory based on its current state. This strategy is akin to RL, enabling the controller to autonomously identify its optimal path, hence synthesizing the tasks of planning and control. The tunnel’s dimensions, determined by the nominal and inner gate values, $W_{n}$ and $W_{gate}$, provide a mechanism to intuitively trade off safety against performance. This balance was adjustable in test runs without necessitating re-tuning, indicating the policy’s robustness across varying tunnel widths.
@@
In Figure 5, we show the trajectory from 10 episodes (30 laps) for both MPCC and SMPCC, noting SMPCC’s paths are notably more consistent. We attribute the inconsistency of MPCC to its representation of the contour weight as a multivariate gaussian introduces numerical asymmetry in the presence of mismatch, as the controller forcefully tries to follow the reference path, and acts poorly in the presence of disturbances.
@@
Figure 5: Simulation experiments of MPCC with the proposed SMPCC, both tuned using TuRBO
@@
We introduce the Episode Success Rate (ESR) metric to quantify episodes completed without gate collisions. As shown in Table II, MPCC achieves an ESR of roughly 70% across all simulations, indicating that 30% of the episode evaluations are not completed successfully due to gate collisions, SMPCC approaches an ESR of nearly 100%. Failed episodes are assigned with the lowest attainable reward and do not provide valuable insights for the BO’s surrogate model updates. As such, we consider these as fruitless interactions with the environment.
@@
SMPCC (ours) Nominal TurBO 99.5 100 99.8
@@
TABLE II: BO Exploration Success Rate: percentage of episodes during training which successfully complete all laps without gate collisions.
V-A3 SMPCC with Model Augmentation
Following the approach described in Section IV-B, we augment the nominal dynamics of the MPC $f(x,u)$ with a residual component $g(x.u)$. Since the residual term is a function of $v_{x}$, $v_{y}$, $v_{z}$, $\Omega^{2}$, requiring to adapt our SMPCC scheme to model the motor speeds $\Omega_{i}$, rather than the direct thrusts $f_{i}$. To this end, we model the motor dynamics as a first-order system and define the states and inputs as described in Section IV-B. In this case, increasing the number of states does not have a significant impact on the solver time in practice.
It’s also worth noting that the motor speeds are generally unavailable at runtime. We thus resort to estimating these online given the current state and applied command. These can be easily estimated from the motor model and low level betaflight controller. Note also that including the residual terms in the MPC formulation potentially introduces more numerical instability, as the features are only plausible within a specific range. To prevent this, we impose an additional constraint on the motor speeds and velocity magnitude to prevent the features from estimating beyond the range in which they characterize the dynamics accurately. We find that we are able to further reduce the lap time on all of the environments by approximately 0.1s. We attribute this improvement to a combination of the above: i) respecting the motor dynamics; ii) the actual augmentation.
V-B SMPCC with Domain Randomization
@@
We address the robustness of the control policy against different noise realizations. The goal is to find policies which perform well amid model discrepancies. Such mismatches arise from using e.g. different hardware, leading to slight changes in mass, inertia, or thrust. We adapt the existing training pipeline to account for such noise realizations. For each iteration of the BO loop, we execute 10 different episodes each subjected to a distinct noise realization. We then collect the average reward over the 10 episodes and update the BO surrogate as usual. For fairness in comparison, we maintain the same number of interactions with the environment, which implies that the number of BO iterations is reduced by a factor 10 (from 600 to 60). In practice, reducing the number of BO iterations did not show any major limitations. We compare the obtained policies against the nominal SMPCC and observe: i) slight increase in the lap time; ii) in both cases the consistency of the trajectory is preserved; iii) 10 noise realizations are not sufficient for a truly robust policy. We conclude that the control formulation at hand is inherently robust and is able to adapt to both changes in the platform, as well as sudden changes in the dynamics. We owe this property to the controller’s replanning capability within the constraint boundaries.
V-C SMPCC vs RL
As shown in Table I, our SMPCC strategy attains lap times comparable to the leading RL policies [65]. We refer to [7], for a detailed comparison between Optimal Control vs RL. We discuss the difference between the trajectories of both in real world in Section V-D, and provide some general remarks herein. Despite both policies achieving similar performance, SMPCC ensures safety by design. We also find that it’s possible to fly with the same SMPCC policy through a variety of tracks and achieve competitive lap times without retuning, indicating good generalization, while this is generally not the case for RL policies. It’s worth noting that these advantages come at the price of more engineering efforts than an end to end learning framework. In particular, the need to account explicitely for delay compensation, solver infeasibility, adapting the MPC outputs to low level controller commands, alongside managing real-time computational solver limitations.
Figure 6: Real-world flight comparison of the MPCC controller with the Tunnel MPC (TMPC) controller proposed in this work. Even when both controllers have been tuned with TuRBO, we show that the TMPC is able to achieve faster speeds, more consistency than its MPCC counterpart and 100% success rate with no gate crashes.
V-D Real World
@@
We test our approach in the real world with a high-performance, racing drone with a high thrust-to-weight ratio (TWR). We use the Agilicious platform [63] for the real world deployment, as detailed in [7] under the designation 4s drone. The control framework was deployed on ACADOS, using SQP_RTI for real-time computation, with the control loop runs at 100Hz. We use a horizon length of $N=20$ at a prediction rate of 25Hz, resulting in a prediction span of $T=0.8s$.
@@
Our method runs on an offboard desktop computer equipped with an Intel(R) Core(TM) i7-8565U CPU @ 1.80GHz. A Radix FC board that contains the Betaflight [^2] https://www.betaflight.com firmware is used as a low level controller. This low-level controller takes as inputs body rates and collective thrusts. An RF bridge is employed to transmit commands to the drone.
For state estimation, we use a VICON system with 36 cameras that provide the platform with millimeter accuracy measurements of position and orientation at a rate of 400 Hz.
@@
We compute the lap times and success rates as done for the simulation experiments, i.e. averaging over 10 trials (3 laps per trial). In Fig. 6 we show the comparison between the real world runs of the baseline MPCC, SMPCC (ours) and the best RL. Both MPCC and SMPCC are tuned with TuRBO. SMPCC consistently satisfies the tunnel bounds, while achieving higher speeds than the MPCC. Our policy did not crash a single time during the 10 trials, being the first approach to achieve a success rate of 100% in real world. The increase in robustness comes without a compromise in performance, as our approach achieves similar lap times to the best RL controller. SMPCC plans a trajectory very close to the edge of the gates, while in RL the deviation from the gate center is penalized in the reward, hence encouraging the drone to fly closer to the gate center. The Split-S maneuver at $x=-4.3m$, stands out as a critical test of each approach’s characteristics, as it’s performed differently in all three cases. This is the most challenging maneuver of the track and has a significant influence on the overall lap time.
Figure 7: Real world thrust and velocity profiles for MPCC, SMPCC and RL. We note how the proposed approach is able to keep the platform extremely close to the limit of handling.
@@
Figure 7 further explores the differences in throttle level, representing normalized commanded collective thrust, and velocity magnitude among the approaches. While velocity profiles were similar, the commanded throttle values differ significantly between the MPC approaches and RL. This is because in MPC, the solver output needs to be translated to equivalent collective thrust and desired body rate values, which are then commanded to the low level controller. This mapping is non trivial, and needs to account for delay effects. We believe that it has a significant influence on the overall performance and is a major drawback of MPC approaches. On the other hand, RL directly outputs a suitable command for the low level controller.
@@ VI Discussion
This paper introduces SMPCC, a Safe MPCC controller for drone racing that prevents collisions against the gates. We augment the nominal dynamics of the controller with a residual term that is obtained from real-world data and captures unmodeled behaviors such as aerodynamic effects and propeller forces. We tune the controller parameters using TuRBO, a state-of-the-art Bayesian optimization algorithm that efficiently compromises between local and global exploration. Our approach addresses several of the limitations encountered in the baseline MPCC [14]: i) robustness against gate collisions; ii) suboptimality from following a predefined optimal trajectory; iii) tuning effort incurred from tuning a high dimensional parameter space. We benchmark our approach against the baseline MPCC and the best-performing RL controller in both simulation and real-world experiments and show that we are on par in terms of lap times while attaining safety against collisions. Our approach is the first to achieve 100% success rate in real-world experiments.
Nonetheless, SMPCC also presents several limitations. First, there is a need for a centerline that passes through the gate centers and is used as the basis to construct the safety set. While there is no computational effort incurred from it, as it can be almost any smooth spline, it does require some manual effort to be determined. Additionally, unlike RL architectures, our approach restricts the platform to operate within a predefined safety set. This limitation hinders its ability to freely explore and develop new behaviors necessary for executing complex maneuvers, such as loops or ladders, which are commonly seen in drone racing environments.
It remains for future work to understand how the choice of the centerline influences the solution and find a method that generalizes to all track configurations. Second, our approach takes significantly longer to train than RL, as the environment cannot be parallelized on the GPU. This is because the solver calls cannot be batched into tensor operations, as done with MLP based policies. Third, the performance of our controller is limited at run time as it requires an optimization problem to be solved at each step, whereas RL methods require a simple forward pass of the neural network.
Thus, we conclude that the main advantage of our approach is that it provides an intuitive lever to trade-off between performance and robustness. By shrinking the width of the tunnel, the quadrotor flies safely through the track. We can then gradually adjust the width to increase performance. Adjustments in the tunnel width can be made directly without further fine-tuning.

### v2

@@ V-A Simulation
At test time, we select the best policy from the training phase as the one with the highest reward, and run 10 episodes (equivalent to 30 laps). The results reported in Table I are based on the average lap time over the 10 episodes evaluated at test time. Additionally, we introduce the Training Success Rate (TSR) to quantify the percentage of training episodes completed successfully without gate collisions (out of 600). We also employ the Success Rate (SR) metric, as introduced in [7], to quantify the percentage of successful episodes during testing (out of 10).
@@ V-A1 TuRBO vs WML
We first investigate the difference between TuRBO [60] and WML [16], and discuss several advantages of TuRBO. We train the baseline MPCC [14] using both variants. Despite both methods achieving comparable maximum rewards, in this section we delve into understanding how each of these explore the parameter space. Figure 2 illustrates two of the tuned parameters - the minimum contour weight $Q_{c}$ and the progress weight $\mu$ over the number of episodes.
Figure 2: Parameter exploration using TuRBO and WML. We plot two of the tuning parameters - $Q_{c}$ and $\mu$ over the number of episodes. TuRBO’s trust regions enhance exploration of the entire parameter space, while WML randomly restarts the exploration every 100 episodes.
To encourage exploration, WML initiates a new trial every 100 iterations with random initialization, aiming to discover new parameter regions. In contrast, TuRBO employs a more strategic approach by selecting trust regions in a similar manner to how it selects BO points, using informed reasoning. An additional advantage of TuRBO is that the trust regions are independent, allowing parallel execution without incurring extra computational cost. Consequently, we advocate for TuRBO as our preferred tuning algorithm for two reasons: i) comprehensive exploration of the entire parameter space; and ii) computational efficiency and scalability, which are crucial for exploring large parameter spaces.
V-A2 MPCC++ vs MPCC
@@
We train both MPCC++ and the baseline MPCC using TuRBO. We make some practical adjustments to the MPCC++ formulation introduced in Section IV-A. Specifically, we relax the hard constraints of the tunnel and impose soft constraints instead. While our MPCC++ optimization problem provides numerical stability in settings that do not require real-time computations, employing real-time solvers does not inherently ensure numerically stable solutions. Even if a solution is found, real-world implementation presents challenges in ensuring the subsequent state respects the predefined constraints. This is particularly true when planning near boundaries, such as tunnel edges, where minor discrepancies due to model mismatches, disturbances, or state-estimation drift can lead to violations of these constraints in the subsequent states. Soft constraints are commonly employed to address such scenarios and can be performed in a systematic way [64]. This has several benefits: i) improves the numerical stability of the solver; ii) handles infeasibilities which arise from model mismatch by allowing minor violations; and iii) maintains the low computational complexity of MPCC. This adaptation does not have any practical implications in terms of performance or safety. The soft constraint is implemented using a barrier function $p(h(x))$ to embed the constraint $h(x)\geq 0$ into the cost function:
$p(h(x))=\log\left(1+\exp\left(-\alpha h(x)\right)\right)$ (17)
@@
where the penalty slope is set to $\alpha=100$. Our results, as detailed in Table I, demonstrate MPCC++’s superiority over MPCC in terms of lap times and success rates. Due to the track constraint, MPCC++ prevents gate crashes consistently without compromising the lap time. Allowing the controller to independently navigate within the tunnel, instead of following a predefined path, gives it greater flexibility to identify the best trajectory based on its current state. The tunnel dimensions, $W_{n}$ and $W_{gate}$, provide a mechanism to intuitively trade off between performance and safety. This balance was adjustable in test runs without necessitating retuning, indicating the policy’s robustness across varying tunnel widths.
@@
In Figure 3, we show the trajectory from 10 episodes (30 laps) for both MPCC and MPCC++, noting MPCC++’s paths are notably more consistent. We attribute the inconsistency of MPCC to its representation of the contour weight as a multivariate Gaussian, since the controller forcefully tries to follow the reference path, and acts poorly in the presence of disturbances.
@@
Figure 3: Simulation experiments of MPCC with the proposed MPCC++, both tuned using TuRBO.
@@
As shown in Table II, MPCC achieves a Training Success Rate (TSR) of roughly 70% across all simulations, indicating that 30% of the episode evaluations are not completed successfully due to gate collisions, while MPCC++ reaches a TSR of nearly 100%. Failed episodes are assigned with the lowest attainable reward and do not provide valuable insights for the BO’s surrogate model updates. As such, we consider these as fruitless interactions with the environment.
@@
MPCC++ (ours) Nominal TuRBO 99.5 100 99.8
@@
TABLE II: BO Training Success Rate (TSR): percentage of episodes during training completed successfully without gate collisions.
V-A3 MPCC++ with Model Augmentation
Following the approach described in Section IV-B, we augment the nominal dynamics of the MPCC $\bm{f(x,u)}$ with a residual component $\bm{g(x,u)}$. Incorporating polynomial features to augment the dynamics requires knowledge of the motor speeds, which were not present in the original MPCC framework. To address this, we introduce several adaptations.
First, we include a first-order motor model into the dynamics to estimate the motor speeds in real-time:
$\dot{\bm{\Omega}}=\frac{1}{\tau_{mot}}(\bm{\Omega_{des}}-\bm{\Omega})$ (18)
where $\bm{\Omega_{des}}$ and $\bm{\Omega}$ denote the desired and the actual motor speeds, respectively. Second, we require an accurate estimate of these speeds to initialize the controller at each timestep. Additionally, we reformulate the optimal control problem to use motor speeds as inputs instead of single rotor thrusts. These adaptations are essential for leveraging data-driven dynamic features within a real-time optimization framework, setting our approach apart from prior works like [6], which used augmented dynamics for environment simulation in RL.
As shown in Table I, we are able to further reduce the lap time on all of the environments by approximately 0.1s. We attribute this improvement to a combination of the above: i) respecting the motor dynamics; and ii) the actual augmentation.
V-A4 MPCC++ with Domain Randomization
@@
We address the robustness of the control policy against different noise realizations. The goal is to find policies which perform well amid model discrepancies. Such mismatches arise from using e.g. different hardware, leading to slight changes in mass, inertia, or thrust. We adapt the existing training pipeline to account for such noise realizations. For each BO iteration , we execute 10 different episodes each subjected to a distinct noise realization. We then collect the average reward over the 10 episodes and update the BO surrogate as usual. For fairness in comparison, we maintain the same budget of episode evaluations, which implies that the number of BO iterations is reduced by a factor 10 (from 600 to 60). In practice, reducing the number of BO iterations did not show any major limitations. We compare the obtained policies against the nominal MPCC++ and observe: i) slight increase in the lap time; ii) the consistency of the trajectory is preserved; and iii) 10 noise realizations are not sufficient for a truly robust policy. We conclude that the control formulation at hand is able to adapt to both changes in the platform, as well as sudden changes in the dynamics. We owe this property to the controller’s replanning capability within the constraint boundaries.
V-A5 MPCC++ vs RL
In this section, we compare our MPCC++ method against the RL policy from [7]. We refer to the cited work for a detailed comparison between optimal control and RL. As shown in Table I, both MPCC++ and RL achieve similar lap times. However, MPCC++ incorporates explicit safety considerations into its design, enabling it to optimize trajectories by strategically taking shortcuts through gate corners. In practice, MPCC++ employs a low contour weight, $Q_{c}$, which, due to the explicit safety constraints in its design, allows the solver to focus on planning the optimal trajectory without excessively penalizing the contour error. In contrast, RL penalizes the gate passing distances in its reward function, which encourages the drone to navigate through the gate center. A detailed comparison of the trajectories is provided in Section V-B.
While MPCC++ incorporates specific safety measures into its design, it requires considerable engineering effort compared to an end-to-end learning framework like RL. This includes managing real-time computational constraints of the solver, adapting MPC outputs to low-level controller commands, addressing solver infeasibilities, and explicitly compensating for delays.
Figure 4: Real world flight trajectories on the Split-S track for the baseline MPCC [14], MPCC++ and the best-performing RL policy [7]. Both MPCC and MPCC++ were tuned using TuRBO.
V-B Real World
@@
We test our approach in the real world using a high-performance racing drone with a high thrust-to-weight ratio (TWR). We use the Agilicious platform [63], as introduced in [7] under the designation 4s drone. The control framework was deployed on ACADOS [65] using SQP_RTI for real-time computation, with the control loop running at 100Hz. We use a horizon length of $N=20$ at a prediction rate of 25Hz, resulting in a prediction span of $T=0.8s$.
@@
Our method runs on an offboard desktop computer equipped with an Intel(R) Core(TM) i7-8565U CPU @ 1.80GHz. A Radix FC board equipped with the Betaflight [^2] https://www.betaflight.com firmware is used as the low-level controller, which takes as inputs desired body rates and collective thrusts. An RF bridge is employed to transmit commands to the drone. For state estimation, we use a VICON system with 36 cameras that provide the platform with millimeter accuracy measurements of position and orientation at a rate of 400 Hz.
@@
We select the best policy obtained in the residual simulator and deploy it in the real world. We compute the lap times and success rates as done for the simulation experiments, i.e. averaging over 10 episodes. In Figure 4, we show the comparison between the baseline MPCC, MPCC++ and the best-performing RL policy. Both MPCC and MPCC++ are tuned with TuRBO. MPCC++ consistently satisfies the tunnel bounds, while achieving higher speeds than MPCC. Our policy did not crash a single time, being the first approach to achieve a 100% success rate in real-world experiments. The improvement in safety comes without a compromise in performance, as our approach achieves similar lap times to the best-performing RL policy. MPCC++ plans a trajectory that takes shortcuts through the gate corners, while in RL the deviation from the gate center is penalized in the reward, hence encouraging the drone to fly closer to the gate centers. The Split-S maneuver at $x=-4.3m$ and $y=-5.6m$, stands out as a critical test of each approach’s characteristics, as it’s performed differently in all three cases. This complex maneuver requires the drone to fly through a higher gate and then immediately descend through a second gate located directly below the first, with both gates sharing the same $x,y$ coordinates. This is the most challenging maneuver of the track, significantly influencing the overall lap time.
Figure 5: Real world thrust and velocity profiles for MPCC, MPCC++ and RL.
@@
Figure 5 illustrates the differences in throttle level, representing normalized commanded collective thrust, and velocity magnitude among the approaches. While velocity profiles were similar, the commanded throttle values differ significantly. This is because for MPCC and MPCC++, the solver output needs to be translated to equivalent desired body rates and collective thrust values to be commanded to the low-level controller. This mapping is non trivial, and needs to account for delay effects. We believe that it has a significant influence on the overall performance and is a major drawback of MPC-based approaches, while RL directly outputs a suitable command.
@@ VI Discussion
This paper introduces MPCC++, a safe MPCC controller for drone racing. We first introduce safety constraints, in the form of a track constraint and a terminal set, which systematically prevent gate collisions and ensure recursive feasibility. We then augment the nominal dynamics of the controller with a residual term that captures unmodeled behaviors such as aerodynamic effects. Finally, we tune the controller parameters using TuRBO, a state-of-the-art BO algorithm that efficiently compromises between local and global exploration. Our approach addresses several of the limitations encountered in the baseline MPCC: i) no systematic prevention against gate collisions; ii) sub-optimality from following a predefined optimal trajectory; and iii) tuning effort incurred from a high dimensional parameter space. We benchmark MPCC++ against the baseline MPCC and the best-performing RL policy in both simulation and real-world experiments and show that we are on par in terms of lap times while attaining safety against gate collisions. Our approach is the first to achieve 100% success rate in real-world experiments.
Nonetheless, MPCC++ also presents several limitations. First, it requires a centerline that passes through the gate centers and is used as the basis to construct the track constraint. While constructing the centerline incurs no computational effort, this approach does introduce certain limitations. Specifically, our approach requires mapping the drone’s current position to a point on the centerline, which is then linked to a progress value. Additionally, constructing the centerline as a spline requires defining the task as a sequence of waypoints in a known order. In contrast, RL architectures often use reward functions (sparse, binary, etc) to specify tasks, allowing for a simpler and more flexible task specification.
Second, our approach takes significantly longer to train than RL, as the environment cannot be parallelized on the GPU. This is because the solver calls cannot be batched into tensor operations, as done with learning-based policies.
Third, while our approach assumes that gate positions are known with high accuracy, future work should explore incorporating gate position uncertainty into our method.
Finally, our approach requires solving an optimization problem online at each step, which may limit its applicability on onboard computers with limited processing capabilities. In contrast, RL methods only require a simple forward pass of the neural network at runtime.
We conclude that the main advantage of MPCC++ is that it provides an intuitive lever to trade off between performance and safety. By shrinking the width of the tunnel, the quadrotor flies safely through the track. We can then gradually adjust the width to increase performance. Adjustments in the tunnel width can be made directly without the need of further re-tuning.